# Automated Testing of Spatially-Dependent Environmental Hypotheses through Active Transfer Learning

### 03 Mar 2024 v2

@@ Abstract
The efficient collection of samples is an important factor in outdoor information gathering applications on account of high sampling costs such as time, energy, and potential destruction to the environment. Utilization of available a-priori data can be a powerful tool for increasing efficiency. However, the relationships of this data with the quantity of interest are often not known ahead of time, limiting the ability to leverage this knowledge for improved planning efficiency. To this end, this work combines transfer learning and active learning through a Multi-Task Gaussian Process and an information-based objective function. Through this combination it can explore the space of hypothetical inter-quantity relationships and evaluate these hypotheses in real-time, allowing this new knowledge to be immediately exploited for future plans. The performance of the proposed method is evaluated against synthetic data and is shown to evaluate multiple hypotheses correctly. Its effectiveness is also demonstrated on real datasets. The technique is able to identify and leverage hypotheses which show a medium or strong correlation to reduce prediction error by a factor of 1.5–6 within the first 5 samples, and poor hypotheses are quickly identified and rejected, having no adverse effect on planning after $\sim$3 samples.
@@ I Introduction
Additionally, correlations often exist between the values of different quantities. One example is the relationship between Nitrogen, soil moisture, and temperature [3]. These relationships can be exploited through the technique of transfer learning. In the context of this work, transfer learning improves predictions of a QOI by using learned patterns in related quantities. This can also be realized through GP through a “multi-task” covariance function that accommodates data from multiple sources.
@@
One additional and significant aspect of dealing with sparse data from multiple quantities is that samples from the different quantities may not line up in space, which is called non-collocated data. This aspect makes calculating correlations directly between quantities impossible. An intelligent use of GP can handle this too by exploiting their property to interpolate values continuously through space. In effect, a spatially-dependent correlation is calculated.
@@ III-A Model Learning
Predictions and uncertainties across the search region are modeled using a GP, which is defined by its mean and covariance functions. As is common, we use a zero mean [2]. For the base of the covariance function, we use the squared exponential. Its simplicity leads to easier derivation and lower computational cost, and it is applicable to spatially-continuous phenomena found in nature. It takes the form
$cov(x_{i},x_{j}|\sigma_{f}^{2},\sigma_{l}^{2})=\sigma_{f}^{2}\exp\left(-\frac{||x_{i}-x_{j}||^{2}}{2\sigma_{l}^{2}}\right),$ (1)
where $x_{i}$ and $x_{j}$ are sample locations, and $\sigma_{f}$ and $\sigma_{l}$ act as the signal standard deviation and characteristic length-scale, respectively. This is a common single-quantity covariance function. It causes locations closer in space to have greater covariance, which means their sample values are assumed to be closer.
@@
$cov(x_{i,u},x_{j,v}|A,\sigma_{l}^{2})=A_{uv}\exp\left(-\frac{||x_{i,u}-x_{j,v}||^{2}}{2\sigma_{l}^{2}}\right),$ (2)
where $u$ and $v$ are quantity indices, and the matrix element $A_{uv}$ is the covariance between two quantities. The covariance function’s two factors serve two roles. The exponential factor serves as a spatial correlation between samples ranging from 00 to $1$. The matrix $A$ acts as a covariance between entire quantities, i.e. a “quantity-covariance”. See Fig. 3 for an abstract idea of the MTGP architecture.
@@
$\begin{split}\log p(y|X,\theta)&=-\frac{1}{2}\left[y^{T}\Sigma^{-1}y+\log|\Sigma|+n\log 2\pi\right]\\ \Sigma&=K(X,X|\theta)+\sigma_{n}^{2}I,\end{split}$ (4)
where $K(X,X|\theta)$ is the covariance matrix between all training data produced using (2) and $\sigma_{n}^{2}$ is the hyperparameter for the noise variance.
@@ IV Results and Discussion
    Choose next sample location
@@
Simulations were run on all possible combinations of prior data from none to all three classes, making eight tests in all. This was repeated over $18$ runs of randomized map data from which average performance was calculated. This was enough runs to get statistically significant differences between hypothesis groups (p < 0.001). A value for $\alpha$ of $500$ was found to be a good balance with slightly more global exploration, as recommended in [10].
@@ IV-A Hypothesis Evaluations
The algorithm had no problem in correctly evaluating the multiple hypotheses simultaneously. The three classes of prior data (H,M,L) all resulted in consistently distinct hypothesis scores from each other, no matter what the combination. The specific combinations did have slight effects on the scores, especially during the first $5$ or so samples.
Figure 6: Mean prediction errors averaged over all runs. Each series is a different combination of High, Medium, and Low hypotheses. A division into three performance groups based on best available hypothesis is visible after $3$ samples. The value not shown from the L series is $1.22$.
H-priors were assigned perfect scores after the first $2$ samples, or after the first $6$ when combined with a L-prior. At the beginning there is insufficient data to robustly and accurately accept or reject hypotheses, and when two hypotheses widely differ in their predictions, happening to favor the poor one will lead to rejecting the good one. This isn’t a possibility with only hypotheses that agree.
When not combined with a H-prior, M-priors started with scores of almost $1$ and eventually settled at a more reasonable $0.65$. When H-priors were present, they more quickly settled to around $0.5$.
@@ IV-B Accuracy to ground truth
Fig. 6 shows average prediction errors for all $8$ prior data combinations. The algorithm adapted its use of the prior data, and by the second or third sample the errors settled into three distinct groups. These groups are defined by the highest-dependence prior that they include.
All the runs with a H-prior had the lowest errors and maintained that gap for the rest of the $30$ samples. Inclusion of lower-dependence priors slightly degraded performance at the start, but by sample $10$ they were all effectively equal.
The ML-priors run performed just as well as the M-prior run, and the L-prior run performed as well as when including no priors. Low-dependence priors had the greatest effect during the first couple samples — before enough data was collected to discount them. Overall, each run performed no worse than its best hypothesis and incorrect hypotheses were quickly disregarded. Adding moderate to good hypotheses reduced prediction error and thus the number of samples needed to characterize the QOI.
The principal benefits occur in the short to mid-term. Eventually, enough data of the QOI will be gathered so that prior hypotheses no longer make a difference. This can be seen for runs with only a M-prior or worse: after around $16$ samples, they all converged. It is expected that the same thing will occur for all the runs after many more samples. In other words, the QOI will eventually be completely known, whether or not a high-quality hypothesis was used to start.
IV-C Real-World Example
We further demonstrate the usefulness of these methods on two real-world datasets with real environmental relationships. This is to show the main components of the algorithm functioning — the scale and specific quantities can change per outdoor robotic application.
The data were taken from satellite imagery provided by NASA Earth Observations [16]. The two regions are the whole of Australia (Fig. 7) and a patch of land between New South Wales and Queensland, Australia (Fig. 10). The QOI was chosen to be vegetation and $25$ dispersed samples were extracted from each of the other quantities for prior data. The algorithm was set to characterize vegetation across the map during a run of $30$ samples.
@@ IV-C Real-World Example
Figure 8: Estimated correlations to Vegetation over a run of $30$ samples for Australia. It is predicted that rainfall and elevation are highly correlated while ground temperature less so.
Figure 9: Mean absolute prediction errors over a run of $30$ samples for Australia. Using priors reduces errors by more than $50\%$ after $3$ samples as compared to no priors.
For the full Australia dataset, an uncertain period of about $5$ samples was followed by consistent hypothesis evaluations for the rest of the run. Elevation was found to be highly negatively correlated, ground temperature partially negatively correlated, and rainfall highly positively correlated (Fig. 8). Using these good hypotheses improved the prediction accuracy after $3$ samples by $50\%$ and maintained this throughout the rest of the run (Fig. 9).
Since the algorithm calculates a spatially-dependent correlation, a note on length-scale is important here. The amount of correlation in the data between two quantities changes depending on the length-scale of the trends examined. Our algorithm models this by the length-scale hyperparameter in the covariance function (1), and its value is mainly determined by the distance between samples and the variation in their values. Our algorithm focuses on best global fit, so it favors more space between samples and hence a larger, more global length-scale. With that in mind, there may still be local areas that starkly differ from the global trend. For example, the negative correlation found between vegetation and elevation does not hold in the mountainous regions near the southeast coast. Development of strategies to better capture both local and global trends, i.e. correlations on multiple length scales, is planned for future work.
For the local dataset, the method established correlation estimates by around $7$ samples, and they remained mostly stable over the course of the run. The results were partial positive correlations with elevation and rainfall and no correlation with ground temperature (Fig. 11). Mean prediction errors showed approximately equal performance when using priors compared with using no priors (Fig. 12).
@@ V Conclusion
This paper presented a method to address two related tasks in outdoor applications: evaluating spatially-dependent environmental hypotheses and reducing the number of samples to characterize a QOI. The method simultaneously evaluates the quality of multiple hypotheses about linear dependence between quantities after every sample. Furthermore, it adapts its learning to the quality of each hypothesis: good hypotheses increase the amount of information gained from each sample and poor hypotheses have no effect after a few samples. Hypotheses identified as high quality can more confidently be used in further outdoor applications.
@@
Figure 10: Real data maps of vegetation, elevation, ground temperature, and rainfall in a section of east Australia and the sparse samples used. Values are averages over a year and have been filtered and normalized within each map.
Figure 11: Estimated correlations to Vegetation over a run of $30$ samples. It is predicted that rainfall and elevation are partly correlated while ground temperature is not.
Figure 12: Mean absolute prediction errors over a run of $30$ samples. Errors approximately match each other over the course of the runs.

### 07 Mar 2024 v3

@@ Abstract
The efficient collection of samples is an important factor in outdoor information gathering applications on account of high sampling costs such as time, energy, and potential destruction to the environment. Utilization of available a-priori data can be a powerful tool for increasing efficiency. However, the relationships of this data with the quantity of interest are often not known ahead of time, limiting the ability to leverage this knowledge for improved planning efficiency. To this end, this work combines transfer learning and active learning through a Multi-Task Gaussian Process and an information-based objective function. Through this combination it can explore the space of hypothetical inter-quantity relationships and evaluate these hypotheses in real-time, allowing this new knowledge to be immediately exploited for future plans. The performance of the proposed method is evaluated against synthetic data and is shown to evaluate multiple hypotheses correctly. Its effectiveness is also demonstrated on real datasets. The technique is able to identify and leverage hypotheses which show a medium or strong correlation to reduce prediction error by a factor of 1.4–3.4 within the first 7 samples, and poor hypotheses are quickly identified and rejected eventually having no adverse effect.
@@ I Introduction
Additionally, correlations often exist between the values of different quantities. One example is the relationship between Nitrogen, soil moisture, and temperature [3]. These relationships can be exploited through the technique of transfer learning. Though transfer learning is often applied to image and text classification [4, 5], the concept is also useful in domains such as this one. In the context of this work, transfer learning improves predictions of a QOI by using learned patterns in related quantities. This can also be realized through GP through a “multi-task” covariance function that accommodates data from multiple sources.
@@
One additional and significant aspect of dealing with sparse data from multiple quantities is that samples from the different quantities may not line up in space, which is known as non-collocated data. This aspect makes calculating correlations directly between quantities impossible. An intelligent use of GP can handle this too by exploiting their property to interpolate values continuously through space. In effect, a spatially-dependent correlation is calculated.
@@ III-A Model Learning
Predictions and uncertainties across the search region are modeled using a GP, which is defined by its mean and covariance functions. For the base of the covariance function, we use the squared exponential. Its simplicity leads to easier derivation and lower computational cost, and it is applicable to spatially-continuous phenomena found in nature. It takes the form
$cov(\mathbf{x}_{i},\mathbf{x}_{j}|\sigma_{f}^{2},\sigma_{l}^{2})=\sigma_{f}^{2}\exp\left(-\frac{||\mathbf{x}_{i}-\mathbf{x}_{j}||^{2}}{2\sigma_{l}^{2}}\right),$ (1)
where $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ are sample locations, and $\sigma_{f}$ and $\sigma_{l}$ act as the signal standard deviation and characteristic length-scale, respectively. This is a common single-quantity covariance function. It causes locations closer in space to have greater covariance, which means their sample values are assumed to be closer.
@@
$cov(\mathbf{x}_{i,u},\mathbf{x}_{j,v}|A,\sigma_{l}^{2})=A_{uv}\exp\left(-\frac{||\mathbf{x}_{i,u}-\mathbf{x}_{j,v}||^{2}}{2\sigma_{l}^{2}}\right),$ (2)
where $u$ and $v$ are quantity indices, and the matrix element $A_{uv}$ is the covariance between two quantities. The covariance function’s two factors each serve a role. The exponential factor serves as a spatial correlation between samples ranging from 00 to $1$. The matrix $A$ acts as a covariance between entire quantities, i.e. a “quantity-covariance”.
For the mean function, we define a quantity-wise constant mean. That is, for each quantity, it will return the average of all the sample values of that quantity:
$m(\mathbf{x}_{i,u})=\mathbb{E}\left[y_{.,u}\right]$ (3)
This piece, along with the covariance, is essential to producing hypothesis evaluations as described in the next section. See Fig. 3 for an abstract idea of the Multi-Task Gaussian Process (MTGP) architecture.
@@
$\begin{split}\log p(\mathbf{y}|X,\theta)&=-\frac{1}{2}\left[\tilde{\mathbf{y}}^{T}\Sigma^{-1}\tilde{\mathbf{y}}+\log|\Sigma|+n\log 2\pi\right]\\ \Sigma&=K(X,X|\theta)+\sigma_{n}^{2}I\\ \tilde{\mathbf{y}}&=\mathbf{y}-m(X),\end{split}$ (5)
where $K(X,X|\theta)$ is the covariance matrix between all training data produced using (2), $m(X)$ is the array of means produced using (3), and $\sigma_{n}^{2}$ is the hyperparameter for the noise variance and is also trained.
@@ IV Results and Discussion
    Choose next sample location from the model
@@
Simulations were run on all possible combinations of prior data from none to all three classes, making eight tests in all (e.g. see Fig. 6). This was repeated over $18$ runs of randomized map data from which average performance was calculated. This was enough runs to get statistically significant differences between hypothesis groups (p < $0.001$). A value for $\alpha$ of $100$ was found to be a good balance with slightly more global exploration, as recommended in [12].
On a laptop AMD Ryzen 9 6900HS CPU, the simulation took $0.26$ seconds on average to calculate both the belief model and next location after each sample. When run with all $3$ priors with $25$ samples each, it took $2.1$ seconds on average. This can be reduced to below a second if less model refinement is chosen. These are acceptable times for in-situ prediction and decision-making and make real-time use possible.
Figure 6: Mean prediction errors averaged over all runs. Each series is a different combination of High, Medium, and Low hypotheses. A division into three performance groups based on best available hypothesis is visible after $7$ samples. The value not shown from the ML series is $0.54$.
@@ IV-A Hypothesis Evaluations
The algorithm had no problem in correctly evaluating the multiple hypotheses simultaneously. The three classes of prior data (H,M,L) all resulted in consistently distinct hypothesis scores from each other, no matter what the combination. The specific combinations did have slight effects on the scores, especially during the first $5$ or so samples. For any method to robustly and accurately accept or reject hypotheses, sufficient data must be collected.
H-priors were assigned near-perfect scores after the first $6$ samples and maintained this with high confidence throughout the rest of the run. This trend was consistent through all the possible combinations.
When not combined with a H-prior, M-priors started with higher scores that eventually came down to around $0.83$. When H-priors were present, they more quickly settled to around $0.45$. The existence of a good hypothesis helped to learn the distinction between it and the medium one.
@@ IV-B Accuracy to ground truth
Fig. 6 shows average prediction errors for all $8$ prior data combinations. The algorithm adapted its use of the prior data, and by $7$ samples the errors settled into three major groups. These groups are defined by the highest-dependence prior that they include.
Runs with a H-prior had the lowest errors and maintained that gap for the rest of the $30$ samples. Inclusion of lower-dependence priors slightly degraded performance at the start, but by sample $15$ they were all effectively equal.
The ML-priors run performed similarly to the M-prior run, and the L-prior run performed at least as well as when including no priors. Overall, each run performed no worse than its best hypothesis and incorrect hypotheses were quickly disregarded. Adding moderate to good hypotheses reduced prediction error and thus the number of samples needed to characterize the QOI.
The principal benefits occur in the short to mid-term. Eventually, enough data of the QOI will be gathered so that prior hypotheses no longer make a difference. This can be seen for runs with only a M-prior or worse: after around $17$ samples, they all converged. It is expected that the same thing will occur for all the runs after many more samples. In other words, the QOI will eventually be completely known, whether or not a high-quality hypothesis was used to start.
@@
Figure 8: Estimated correlations to Vegetation over a run of $30$ samples for Australia. It is estimated that rainfall is positively correlated while ground temperature and elevation are negatively.
Figure 9: Mean absolute prediction errors over a run of $30$ samples for Australia. Using priors reduces errors over much of the run as compared to no priors.
IV-C Real-World Examples
We further demonstrate the usefulness of these methods on two real-world datasets with real environmental relationships. This is to show the algorithmic methods of this paper working — the scale and specific quantities can change per outdoor robotic application. Also note that the only thing simulated in these tests is the action of sampling — the methods perform the same on the data regardless of how or when it was obtained. As a plus, having high-resolution offline data allows for ground truth comparisons not possible in the field.
The data were taken from satellite imagery provided by NASA Earth Observations [18]. The two regions are the whole of Australia (Fig. 7) and a 1000x1000 km patch of land between New South Wales and Queensland, Australia (Fig. 10). The QOI was chosen to be vegetation and $25$ dispersed samples were extracted from each of the other quantities for prior data. The algorithm was set to characterize vegetation across the map during a run of $30$ samples.
After only $2$ samples on the full Australia dataset, positive and negative correlations were identified and maintained for the rest of the run. Elevation and ground temperature were found to be highly negatively correlated and rainfall highly positively correlated (Fig. 8). Using these good hypotheses improved the prediction accuracy throughout almost the entire run, including up to $75\%$ better in the beginning (Fig. 9).
Since the algorithm calculates a spatially-dependent correlation, a note on length-scale is important here. The amount of correlation in the data between two quantities changes depending on the length-scale of the trends examined. Our algorithm models this by the length-scale hyperparameter in the covariance function (1), and its value is mainly determined by the distance between samples and the variation in their values. Our algorithm focuses on best global fit, so it favors more space between samples and hence a larger, more global length-scale. With that in mind, there may be local areas that differ from the global trend, especially with real data. For example, the negative correlation found between vegetation and elevation will not hold in the mountainous regions near the southeast coast. Development of strategies to capture both local and global trends, i.e. correlations on multiple length scales, is of interest for future work.
Figure 10: Real data maps of vegetation, elevation, ground temperature, and rainfall in a section of east Australia and the sparse samples used. Values are averages over a year and have been normalized within each map.
For the local dataset, the method established correlation estimates by around $7$ samples, and they remained mostly stable over the course of the run. The results were partial positive correlations with elevation and rainfall and a negative correlation with ground temperature (Fig. 11). Mean prediction errors showed slightly improved performance when using priors compared with using no priors (Fig. 12).
@@ IV-C Real-World Examples
Figure 11: Estimated correlations to Vegetation over a run of $30$ samples. It is predicted that all quantities are partly correlated with vegetation.
Figure 12: Mean absolute prediction errors over a run of $30$ samples. Using priors slightly improves over using no priors, particularly at the end.
@@ V Conclusion
This paper presented a method to address two related tasks in outdoor applications: evaluating spatially-dependent environmental hypotheses and reducing the number of samples to characterize a QOI. The method simultaneously evaluates the quality of multiple hypotheses about linear dependence between quantities after every sample. Furthermore, it adapts its learning to the quality of each hypothesis: good hypotheses increase the amount of information gained from each sample and poor hypotheses lose effect after a few samples. Hypotheses identified as high-quality can more confidently be used in further outdoor applications.